# Language Model Fine-Tuning on Scaled Survey Data for Predicting Distributions of Public Opinions

### v1

@@ Appendix B Experiment Details
[…]
@@ C.1 Effect of Response Distribution Modeling
Table 5 summarizes the results of these approaches. Notably, explicit probability modeling substantially outperforms the one-hot method, demonstrating that simply predicting the single most frequent response fails to capture the opinion diversity present within each subpopulation.
Compared with augment by $N$ (2nd and 3rd column in Table 5), explicit probability modeling also achieves better performance. Importantly, the performance gap exceeds the quantization error introduced by discretizing the response distribution. For instance, when discretizing with a factor of $N$, the quantization error is $\frac{1}{2N}$—approximately 0.01 or 0.005 in the cases shown in Table 5. Moreover, explicit modeling offers the practical benefit of reducing the data volume by a factor of $N$ compared to the augmentation approach, thereby lowering the computational cost of fine-tuning LLMs.
@@
Table 5: Comparison of evaluation performance for three response distribution modeling approaches, with Llama-2-7B as a base model. The last column (Explicit) is identical to the ours presented in Table 1. A model fine-tuned to predict the most probable choice (one-hot) performs the worst, as the model has not learned distributional opinion at fine-tuning phase. A model trained on augmented data (Aug. ($\times$50, $\times$100)), while performing much better than one-hot still underperforms the explicit distribution modeling.
@@ C.2 Post-trained Model
We fine-tune Llama-2-7B-chat to observe the effect of starting from checkpoints that have been instruction-tuned via Reinforcement Learning from Human Feedback (RLHF). Table 6 shows the evaluation performance of a baseline method (Zero-shot prompt (QA)), fine-tuned base model (Llama-2-7B) and fine-tuned chat model (Llama-2-7B-chat). We observe the significant performance improvement, while the baseline method performs worse then the models not instruction-tuned (Table 1). Especially, the performance for SubPOP-Eval of chat model is significantly worse than that of base model. We observe the high WD of the baseline method resulting from the model assigning high probability to a specific token (e.g. ‘A’), being far apart from the human opinion distribution. After fine-tuning the model are able to generate a more distributed probability of answer tokens. This result coincides with the result reported in Moon et al. (2024).
@@ C.3 Generalization to Unseen Subpopulations
As shown in Table 7, we observe a performance improvement across unseen subpopulations. To verify that the performance improvements arise from the fine-tuned model being able to steer towards given subpopulations, we measure intergroup disagreement pattern for the demographic and ideology traits, shown in Figure 10, 11, 12, and 13. We consistently observe across traits that the disagreement pattern of our model resembles that of the human group, while zero-shot prompting with the base model exhibits a pattern completely different from the human group result. This observation shows that our fine-tuned model learns to condition on subpopulation information and also generalizes to subpopulations unseen during fine-tuning.
Table 7: Evaluation performance of our fine-tuned Llama-2-7B model on OpinionQA for subpopulations not included in the fine-tuning dataset SubPOP-Train. For reference, we present a lower bound (human) and the zero-shot prompting (QA). Absolute difference refers to the WD difference between zero-shot prompting and ours, and the relative improvement is calculated in a same way as Figure 3.
@@ Appendix D Baseline Details
  • Zero-shot prompting: Three prompt styles—QA, BIO, and PORTRAY—are introduced in Santurkar et al. (2023) to integrate group information into prompts. These prompts are then combined with survey questions to construct inputs for LLM. Then, the first-token log-probability from LLM is measured to calculate the model’s response distribution over options. In our baseline (and also in fine-tuning experiments) we focus on the QA steering format. Examples of this prompting method are shown in Figure 14.
@@
    $\mathcal{D}_{g}^{(H)}$ =$=\left\{\frac{1}{|Q|}\sum_{q\in Q}WD(\pi_{gq}^{(H)},\pi_{gq}^{(r)})\right\}_{r=1}^{R}.$

### v2

@@ Appendix B Experiment Details
$\mathcal{WD}(p_{H},p_{\theta})$ $=\int_{-\infty}^{+\infty}\left|F_{p_{H}}(x)-F_{p_{\theta}}(x)\right|\,dx,$
@@ C.1 Effect of Response Distribution Modeling
Table˜5 summarizes the results of these approaches. Notably, explicit probability modeling substantially outperforms the one-hot method, demonstrating that simply predicting the single most frequent response fails to capture the opinion diversity present within each subpopulation.
Compared with augment by NN (2nd and 3rd column in Table˜5), explicit probability modeling also achieves better performance. Importantly, the performance gap exceeds the quantization error introduced by discretizing the response distribution. For instance, when discretizing with a factor of NN, the quantization error is $\frac{1}{2N}$—approximately 0.01 or 0.005 in the cases shown in Table˜5. Moreover, explicit modeling offers the practical benefit of reducing the data volume by a factor of NN compared to the augmentation approach, thereby lowering the computational cost of fine-tuning LLMs.
@@
Table 5: Comparison of evaluation performance for three response distribution modeling approaches, with Llama-2-7B as a base model. The last column (Explicit) is identical to the ours presented in Table˜1. A model fine-tuned to predict the most probable choice (one-hot) performs the worst, as the model has not learned distributional opinion at fine-tuning phase. A model trained on augmented data (Aug. ($\times$50, $\times$100)), while performing much better than one-hot still underperforms the explicit distribution modeling.
@@ C.2 Post-trained Model
We fine-tune Llama-2-7B-chat to observe the effect of starting from checkpoints that have been instruction-tuned via Reinforcement Learning from Human Feedback (RLHF). Table˜6 shows the evaluation performance of a baseline method (Zero-shot prompt (QA)), fine-tuned base model (Llama-2-7B) and fine-tuned chat model (Llama-2-7B-chat). We observe the significant performance improvement, while the baseline method performs worse then the models not instruction-tuned (Table˜1). Especially, the performance for SubPOP-Eval of chat model is significantly worse than that of base model. We observe the high WD of the baseline method resulting from the model assigning high probability to a specific token (e.g. ‘A’), being far apart from the human opinion distribution. After fine-tuning the model are able to generate a more distributed probability of answer tokens. This result coincides with the result reported in Moon et al. (2024).
@@ C.3 Generalization to Unseen Subpopulations
As shown in Table˜7, we observe a performance improvement across unseen subpopulations. To verify that the performance improvements arise from the fine-tuned model being able to steer towards given subpopulations, we measure intergroup disagreement pattern for the demographic and ideology traits, shown in Figure˜10, 11, 12, and 13. We consistently observe across traits that the disagreement pattern of our model resembles that of the human group, while zero-shot prompting with the base model exhibits a pattern completely different from the human group result. This observation shows that our fine-tuned model learns to condition on subpopulation information and also generalizes to subpopulations unseen during fine-tuning.
Table 7: Evaluation performance of our fine-tuned Llama-2-7B model on OpinionQA for subpopulations not included in the fine-tuning dataset SubPOP-Train. For reference, we present a lower bound (human) and the zero-shot prompting (QA). Absolute difference refers to the WD difference between zero-shot prompting and ours, and the relative improvement is calculated in a same way as Figure˜3.
@@ Appendix D Baseline Details
  • Zero-shot prompting: Three prompt styles—QA, BIO, and PORTRAY—are introduced in Santurkar et al. (2023) to integrate group information into prompts. These prompts are then combined with survey questions to construct inputs for LLM. Then, the first-token log-probability from LLM is measured to calculate the model’s response distribution over options. In our baseline (and also in fine-tuning experiments) we focus on the QA steering format. Examples of this prompting method are shown in Figure˜14.
@@
    $\mathcal{D}_{g}^{(H)}$ $=\left\{\frac{1}{|Q|}\sum_{q\in Q}WD(\pi_{gq}^{(H)},\pi_{gq}^{(r)})\right\}_{r=1}^{R}.$